OpenStudy (anonymous):

simplify cube root of (a-5)^6

OpenStudy (anonymous):

\[\left( x^m \right)^{\frac{ 1 }{ 3 }}=x ^{\frac{ m }{ 3 }}\]

OpenStudy (anonymous):

so it will become (a-5)^2?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw